TheSmartOne (thesmartone):

Calculus 1 Question Find the equations of both tangent lines to the graph of the ellipse x^2/4 + y^2/9 = 1 that passes through the point (0, 4) not on the graph.

TheSmartOne (thesmartone):

Find the equations of both tangent lines to the graph of the ellipse x^2/4 + y^2/9 = 1 that passes through the point (0, 4) not on the graph.

TheSmartOne (thesmartone):

\(\Large \frac{x^2}{4} + \frac{y^2}{9} = 1\)

TheSmartOne (thesmartone):

That equation is the same thing as 9x^2 + 4y^2 = 36 And differentiating that, we get 18x + 8yy' = 0 Solving for y' we get y' = (-9/4) * (x/y)

TheSmartOne (thesmartone):

We know the slope is the same as the derivative (y - 4)/(x - 0) = (-9/4) * (x/y) (y-4)/x = (-9x)/(4y) Cross multiply 4y^2 - 16y = -9x^2 Same thing as 9x^2 + 4y^2 = 16y And plugging in what we know from the beginning 36 = 16y y = 2.25 And I dunno what to do after this >.>

OpenStudy (visarika23):

Couldn't we simply do like this y'=(-9/4)*(x/y) As the tangent passes through (0,4) So we plug in the values and get y'=0 that the line is parallel to x-axis

TheSmartOne (thesmartone):

Lemme try plugging that back in to the ellipse 9x^2 + 4y^2 = 36 9x^2 + 4(2.25)^2 = 36 9x^2 + 20.25 = 36 x^2 = (36 - 20.25)/9 x^2 = 1.75 \(x = \pm \sqrt{1.75}\)

TheSmartOne (thesmartone):

the point (0, 4) is not on the ellipse so the tangent line will not be having a slope of 0

OpenStudy (visarika23):

I know (0,4) is the point through which the tangent passes and y' is the slope of the tangent at point x,y . So can't we put x=0 and y=4 @TheSmartOne

TheSmartOne (thesmartone):

we did, and that's how we got (y - 4)/(x - 0) = (-9/4) * (x/y)